If f(x) = 4X + 2X^2, find f^!(x) from the first principle and hence find the value of a given that f^!(a) = 5


1
Expert's answer
2021-09-14T06:04:59-0400
"f'(x)=\lim\limits_{h\to 0}\dfrac{f(x+h)-f(x)}{h}"


"=\lim\limits_{h\to 0}\dfrac{4(x+h)+2(x+h)^2-4x-2x^2}{h}"

"=\lim\limits_{h\to 0}\dfrac{4x+4h+2x^2+4xh+2h^2-4x-2x^2}{h}"

"=\lim\limits_{h\to 0}\dfrac{4h+4xh+2h^2}{h}=\lim\limits_{h\to 0}(4+4x+2h)"


"=4+4x+0=4+4x"

"f'(x)=4+4x"


"f'(a)=5"

"4+4a=5"


"4a=1"

"a=\dfrac{1}{4}"